# Unleashing the Power of Neural Collapse: Consistent Supervised-Unsupervised Alignment for Generalized Category Discovery

### v1

@@ 4.2 ETF Alignment
Pre-assigned ETF Prototype. We construct a Simplex Equiangular Tight Frame (ETF) as pre-assigned classifier prototypes to enforce a structured feature space and maximize category separability. Given or estimated a number of categories $K$, we can get its categories set $\mathcal{Y}_{t}$. Then we define the optimal ETF prototypes $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$ as Eq.1.
@@
By pre-assigning the ETF prototypes, we provide a stable and optimal alignment structure for both known and novel categories.
@@
Once pseudo-labels $\hat{y}_{i}$ are assigned, we compute the similarity between each sample and its respective cluster center:
@@
where $\mathcal{D}_{k}$ is the set of samples assigned to cluster k.
To align features with the pre-assigned ETF prototype $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$, we select the top $\alpha\%$ of samples with the highest similarity $s_{i}$ within each cluster. These high-confidence samples are pulled toward the corresponding ETF prototype $p_{k}$ using a Dot-Regression Loss[23]:
@@
where $\tilde{D}_{k}=\{x_{i}\mid s_{i}\text{ is in the top }\alpha\%\text{ of }\mathcal{D}_{k}\}$ represents the selected high-confidence samples assigned to cluster $c_{k}$.
Supervised ETF Alignment. For each sample $x_{i}^{l}\in\mathcal{D}^{l}$, we extract its feature embedding $e_{i}^{l}$. The SMC establishes a one-to-one mapping between predicted clusters and ground-truth labels, transforming the original labels $y_{i}^{l}$ into ETF-aligned labels $y_{i}^{\text{{ETF}}}=\phi_{\text{{SCM}}}({y}_{i}^{l})$ (detailed in Section 4.3). Subsequently, each labeled sample is encouraged to align with its assigned ETF prototype:
@@
where $\gamma$ is a weighting coefficient determining the trade-off between unsupervised and supervised ETF alignment.
@@ 4.3 Semantic Consistency Matcher (SCM)
In periodically clustering-based learning, two key challenges arise: 1) Pseudo-label instability across iterations. Clustering fluctuations introduce inconsistencies, disrupting the learning process and degrading feature alignment. 2) Misalignment between the predicted clusters and the true labels in supervised alignment. Supervised ETF alignment may misassign cluster labels, leading to further optimization instability. To mitigate these issues, we propose SCM, which stabilizes pseudo-label assignments across iterations and ensures consistent alignment between predicted clusters and true labels. By enforcing a one-to-one mapping between clustering results from consecutive iterations, SCM reduces clustering-induced fluctuations.
Specifically, SCM aligns pseudo-labels between consecutive clustering iterations through an optimal assignment strategy. Given prediction sets $\{\hat{y}_{1}^{t},\dots,\hat{y}_{N}^{t}\},\quad\forall\hat{y}_{i}^{t}\in\mathcal{Y}_{t}$ for the current iteration and $\{\hat{y}_{1}^{t-1},\dots,\hat{y}_{N}^{t-1}\},\quad\forall\hat{y}_{i}^{t-1}\in\mathcal{Y}_{t-1}$ for the previous iteration, where $\mathcal{Y}_{t}$ and $\mathcal{Y}_{t-1}$ represent the corresponding label sets. Then we seek an optimal permutation $\sigma^{*}$ that maximizes label consistency:
@@ 4.4 Comparative Representation Learning
The unsupervised loss encourages similarity between embeddings of the same sample under different augmentations while distinguishing them from other samples [7]:
@@
where $H(i)$ represents the set of samples sharing the same label as $i$. The total representation learning objective is:
$\mathcal{L}_{\text{REP}}=(1-\lambda)\mathcal{L}_{\text{REP}}^{u}+\lambda\mathcal{L}_{\text{REP}}^{s}.$ (12)
@@ 4.5 Final Objective
where $\beta$ controls the relative contribution of ETF alignment loss and representation learning loss.
@@
(a) Clustering with the ground-truth number of classes $K$ given
GCD (CVPR 2022) 51.3 56.6 48.7 39.0 57.6 29.9 45.0 41.1 46.9 35.4 51.0 27.0 73.0 76.2 66.5 74.1 89.8 66.3
DCCL (CVPR 2023) 63.5 60.8 64.9 43.1 55.7 36.2 - - - - - - 75.3 76.8 70.2 80.5 90.5 76.2
PromptCAL (CVPR 2023) 62.9 64.4 62.1 50.2 70.1 40.6 52.2 52.2 52.3 37.0 52.0 28.9 81.2 84.2 75.3 83.1 92.7 78.3
GPC (ICCV 2023) 55.4 58.2 53.1 42.8 59.2 32.8 46.3 42.5 47.9 - - - 77.9 85.0 63.0 76.9 94.3 71.0
SimGCD (ICCV 2023) 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8 44.0 58.0 36.4 80.1 81.2 77.8 83.0 93.1 77.9
PIM (CVPR 2023) 62.7 75.7 56.2 43.1 66.9 31.6 - - - 42.3 56.1 34.8 78.3 84.2 66.5 83.1 95.3 77.0
SelEx (ECCV2024) 73.6 75.3 72.8 58.5 75.6 50.3 57.1 64.7 53.3 - - - 82.3 85.3 76.3 83.1 93.6 77.8
CMS (CVPR 2024) 68.2 76.5 64.0 56.9 76.1 47.6 56.0 63.4 52.3 36.4 54.9 26.4 82.3 85.7 75.5 84.7 95.6 79.2
SPT (ICLR 2024) 65.8 68.8 65.1 59.0 79.2 49.3 59.3 61.8 58.1 43.4 58.7 35.2 81.3 84.3 75.6 85.4 93.2 81.4
@@
(b) Clustering without the ground-truth number of classes $K$ given
GCD (CVPR 2022) 51.1 56.4 48.4 39.1 58.6 29.7 - - - 37.2 51.7 29.4 70.8 77.6 57.0 77.9 91.1 71.3
GPC (ICCV 2023) 52.0 55.5 47.5 38.2 58.9 27.4 43.3 40.7 44.8 36.5 51.7 27.9 75.4 84.6 60.1 75.3 93.4 66.7
PIM (CVPR 2023) 62.0 75.7 55.1 42.4 65.3 31.3 - - - 42.0 55.5 34.7 75.6 81.6 63.6 83.0 95.3 76.9
CMS (CVPR 2024) 64.4 68.2 62.2 51.7 68.9 43.4 55.2 60.6 52.4 37.4 56.5 27.1 79.6 83.2 72.3 81.3 95.6 74.2
@@ 5.1 Experimental Setup
Datasets. Our evaluation spans six image classification benchmarks, including both generic and fine-grained datasets. For generic datasets, we use CIFAR-100 [28] and ImageNet-100 [29, 30]. For fine-grained datasets, we evaluate on CUB-200 [31], Stanford Cars [32], FGVC Aircraft [33], and Herbarium19 [34]. To separate classes into known and novel categories, we follow the SSB split protocol [35] for the fine-grained datasets. For CIFAR-100 and ImageNet-100, we perform a random class split using a fixed seed, consistent with previous studies [36]. More details regarding the class splits can be found in the Appendix.
Implementation Details. Our method utilizes the pre-trained DINO ViT-B/16 [37, 38] as the backbone network. We fine-tune the final layer of the image encoder along with a projection head, following prior work [3, 7]. The projection head consists of an MLP with a 768-dimensional input, a 2048-dimensional hidden layer, and a 768-dimensional output, followed by a GeLU activation function [39]. The learning rate is set to 0.1. Other hyperparameters, including batch size, temperature $\tau_{s}$, weight decay $\lambda$, and the number of augmentations, are set to 128, 0.07, $1e^{-4}$, 0.35, and 2, respectively, in accordance with previous studies [36]. All experiments are conducted on an NVIDIA A6000 GPU.
Method Fine-grained datasets Avg Classification datasets Avg All datasets Avg
@@
(a) Clustering with the ground-truth number of classes K given
@@
Ours 60.3 67.6 56.7 85.5 89.8 81.4 68.7 75.0 64.9
Improv. over SPT +3.4 +0.5 +4.8 +2.2 +1.0 +2.9 +3.0 +0.7 +4.2
(b) Clustering without the ground-truth number of classes K given
@@
Ours 55.5 64.6 50.9 83.1 89.8 77.3 64.7 73.0 59.7
Improv. over CMS +3.3 +1.1 +4.6 +2.6 +0.4 +4.1 +3.1 +0.9 +4.4
Sup ETF Alignment Unsup ETF Alignment CUB Stanford Cars FGVC Aircraft Herbarium 19 CIFAR100 ImageNet100 Avg
All Old New All Old New All Old New All Old New All Old New All Old New All Old New
67.7 75.7 63.9 55.1 74.3 45.8 54.3 57.4 52.8 36.5 55.0 26.5 82.0 85.4 75.2 84.4 94.0 80.8 63.3 73.7 57.5
69.6 75.8 66.5 57.0 75.1 46.7 56.1 62.9 53.2 37.6 56.7 28.1 82.3 85.7 75.7 85.0 95.3 79.3 64.6 75.2 58.3
75.7 71.3 77.8 56.0 74.7 46.9 59.8 54.7 62.3 47.2 60.0 40.3 82.0 84.1 77.8 87.6 94.5 84.1 68.0 73.2 64.9
74.8 76.8 73.8 59.9 77.8 51.2 60.0 57.6 61.2 46.4 58.4 40.7 82.7 85.5 77.3 88.4 94.1 85.5 68.7 74.9 64.9
Improv. over baseline +7.1 +0.8 +9.8 +4.8 +3.5 +5.4 +5.7 +0.2 +8.4 +9.9 +3.4 +14.2 +0.8 +0.1 +2.1 +3.9 +0.1 +4.7 +5.3 +1.2 +7.4
Table 3: Performance comparison across different datasets with Supervised and Unsupervised ETF Alignment
5.2 Comparison with State of the Art Methods
We evaluate our NC-GCD framework on several fine-grained and generic classification benchmarks to demonstrate its efficacy in GCD. Tab. 1 presents a comparison between our method and the SOTA methods [3, 4, 5, 10, 11, 6, 8, 7, 40], evaluated on both fine-grained and generic datasets with and without access to the ground-truth (GT) number of classes K for clustering. For the case without access to GT K, we estimate the K through clustering, which is then used to align the ETF prototypes. Meanwhile, to comprehensively evaluate the performance of various methods across Fine-Grained Datasets, Generic Classification Datasets, and All datasets, Table 2 presents the average performance.
Fine-Grained Datasets. On fine-grained benchmarks such as CUB-200, Stanford Cars, FGVC Aircraft, and Herbarium19, NC-GCD consistently outperforms previous methods, whether or not the ground-truth K is provided. As shown in Table 2 (a), when the GT K is available, NC-GCD achieves an average all-class accuracy of 60.3%, surpassing SPT by 3.4% and improving novel class accuracy over SPT by 4.8%. Even without GT K, as shown in Table 2 (b), NC-GCD remains robust, surpassing CMS by 3.3% in all-class accuracy. Notably, on CUB-200, NC-GCD outperforms CMS by 5.9% in all-class accuracy and 7.2% in novel class accuracy, demonstrating strong generalization even without the true number of categories. Fine-grained datasets present higher category confusion due to minimal intra-class variation, and our method effectively mitigates this and enhances class separability.
Generic Classification Datasets. NC-GCD achieves strong performance on standard classification datasets such as CIFAR-100 and ImageNet-100, as shown in Table 2. With known K, NC-GCD attains an all-class accuracy of 88.7% on CIFAR-100 and 94.1% on ImageNet-100, outperforming CMS and SPT. Notably, novel class accuracy improves by 3.0% on CIFAR-100 and 2.9% on ImageNet-100. Even when K must be estimated, NC-GCD maintains its competitive edge, achieving 55.5% all-class accuracy, surpassing CMS by 3.3%, demonstrating its adaptability in generic classification GCD scenarios.
Across both fine-grained and generic datasets, NC-GCD consistently achieves strong performance. As shown in Table 2 (a), with GT K, NC-GCD reaches the highest all-class accuracy of 68.7%, surpassing SPT by 3.0% and improving novel class accuracy by 4.2%. Even without GT K, NC-GCD remains robust, surpassing CMS by 3.1% in all-class accuracy and 4.4% in novel class accuracy (Table 2 (b)). By pre-assigning ETF prototypes, NC-GCD effectively improves class separability, mitigating category confusion and ensuring balanced learning across old and novel categories. These results highlight the robustness and versatility of NC-GCD for real-world GCD applications.
@@ 5.3 Ablation Study
Unsupervised ETF Alignment: This component yields the most significant improvements, especially in fine-grained datasets. As shown in Table 3, enabling unsupervised ETF alignment alone leads to substantial gains in novel class accuracy. The model effectively learns robust representations, particularly for novel categories, where differentiation between visually similar classes is crucial.
Supervised ETF Alignment: This component stabilizes and preserves old classes’ accuracy. By aligning labeled features with their corresponding ETF prototypes, it ensures the retention of learned knowledge while maintaining high performance on known categories. Though its contribution to novel class accuracy is smaller than that of unsupervised alignment, it plays a crucial role in preventing forgetting and maintaining model stability.
Synergistic Combination: The full model, integrating both alignments, achieves the highest overall accuracy, with an average all-class accuracy of 68.7%, significantly surpassing the baseline and individual variants. This combination enhances novel class discovery while preserving old class performance. Specifically, the average novel class accuracy increases by 7.4%, and the overall accuracy improves by 5.3%. The results highlight the synergy between supervised and unsupervised alignment, demonstrating the NC-GCD framework’s effectiveness in balancing learning across known and novel categories.
@@
Figure 5: Visualization of our method training process on CUB-200.
SCM CUB Herbarium 19 ImageNet100
ALL Old Novel ALL Old Novel ALL Old Novel
70.3 71.0 70.0 42.6 54.3 36.3 84.3 88.9 82.0
75.7 71.3 77.8 47.2 60.0 40.3 87.6 94.5 84.1
Improv. +5.3 +0.4 +7.8 +4.5 +5.6 +3.9 +3.3 +5.6 +2.1
Table 4: Ablation study of our SCM module
The Effectiveness of SCM. To evaluate the impact of the Semantic Consistency Matcher (SCM), we conduct ablation studies on CUB-200, ImageNet100, and Herbarium 19. As shown in Table 4, enabling SCM consistently improves performance on all three datasets. On Herbarium 19, which poses greater challenges due to its large number of categories and high intra-class variance, SCM provides more significant improvements, with gains of 4.5% for all classes and 5.6% for novel classes, demonstrating its effectiveness in complex classification tasks. By mitigating clustering fluctuations, SCM ensures stable pseudo-label assignments. These results highlight its crucial role in stabilizing training, particularly in datasets with large class numbers and high intra-class similarity.
The Influence of the ETF Scope Coefficient $\alpha$. The ETF Scope Coefficient $\alpha$ plays a crucial role in determining the number of samples selected for alignment with the ETF prototype. By controlling the percentage of high-confidence samples, $\alpha$ influences the model’s ability to focus on either old or novel classes. As seen in Table 3, when $\alpha$ is too low, the model may not leverage enough high-confidence samples, leading to suboptimal performance. On the other hand, as $\alpha$ increases, novel class accuracy improves significantly, especially in fine-grained datasets where class distinctions are subtle. However, when $\alpha$ exceeds a certain threshold, the model’s focus on new classes comes at the expense of old class accuracy, highlighting the trade-off between the two. The best balance is achieved at $\alpha=0.8$, where both old and novel class accuracy reach optimal values, demonstrating the importance of selecting an appropriate $\alpha$.
Ablation study of ETF loss coefficient $\beta$ and Parameter analysis will be presented in the Appendix.

### v2

@@ 4.2 ETF Alignment
Pre-assigned ETF Prototype. We construct a Simplex ETF as pre-assigned classifier prototypes to enforce a structured feature space and maximize category separability. Given or estimated the number of categories $K$, we can get its categories set $\mathcal{Y}_{t}$. Then we define the optimal ETF prototypes $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$ as Eq.1.
@@
Here, $q_{i}$ denotes a centered orthonormal basis of a (K1)(K-1)-dimensional space, with unit vectors satisfying $\sum_{i}q_{i}=\mathbf{0}$ and $q_{i}^{\top}q_{j}=\delta_{ij}$. By pre-assigning the ETF prototypes, we provide a stable and optimal alignment structure for both known and novel categories.
@@
Once pseudo-labels $\hat{y}_{i}$ are assigned, we compute the similarity between each sample and its respective cluster center, where $\mathcal{D}_{k}$ is the set of samples assigned to cluster k:
@@
To align features with the pre-assigned ETF prototype $\mathbf{P}=\{p_{1},p_{2},\dots,p_{K}\}$, we select the top $\alpha\%$ of samples with the highest similarity $s_{i}$ within each cluster. These high-confidence samples are pulled toward the corresponding ETF prototype $p_{k}$ using a Dot-Regression LossYang2022b_LearnableClassifier :
@@
[…]
Supervised ETF Alignment. For each sample $x_{i}^{l}\in\mathcal{D}^{l}$, we extract its feature embedding $e_{i}^{l}$. The SCM establishes a one-to-one mapping between predicted clusters and ground-truth labels, transforming the original labels $y_{i}^{l}$ into ETF-aligned labels $y_{i}^{\text{{ETF}}}=\phi_{\text{{SCM}}}({y}_{i}^{l})$ (detailed in Section 4.3). Subsequently, each labeled sample is encouraged to align with its assigned ETF prototype:
@@
where $\gamma$ is a weighting coefficient determining.
@@ 4.3 Semantic Consistency Matcher (SCM)
In periodically clustering-based learning, two key challenges arise: 1) Pseudo-label instability across iterations. Clustering fluctuations introduce inconsistencies, disrupting the learning process and degrading feature alignment. 2) Misalignment between the predicted clusters and the true labels in supervised alignment. Supervised ETF alignment may misassign cluster labels, leading to further optimization instability. To mitigate these issues, we propose the SCM, which stabilizes pseudo-label assignments across iterations and ensures consistent alignment between predicted clusters and true labels. By enforcing a one-to-one mapping between clustering results from consecutive iterations, SCM reduces clustering-induced fluctuations.
Specifically, SCM aligns pseudo-labels between consecutive clustering iterations through an optimal assignment strategy. Given prediction sets $\{\hat{y}_{1}^{t},\dots,\hat{y}_{N}^{t}\},\quad\forall\hat{y}_{i}^{t}\in\mathcal{Y}_{t}$ for the current iteration and $\{\hat{y}_{1}^{t-1},\dots,\hat{y}_{N}^{t-1}\},\quad\forall\hat{y}_{i}^{t-1}\in\mathcal{Y}_{t-1}$ for the previous iteration, where $\mathcal{Y}_{t}$ and $\mathcal{Y}_{t-1}$ represent the corresponding label sets. Then, we seek an optimal permutation $\sigma^{*}$ that maximizes label consistency:
@@ 4.4 Comparative Representation Learning
Following CMSCMS2024 , we adopt the unsupervised loss encourages similarity between embeddings of the same sample while distinguishing them from other samples:
@@
where $H(i)$ represents the samples sharing the same label $i$. The representation learning objective is:
$\mathcal{L}_{\text{REP}}=(1-\lambda)\mathcal{L}_{\text{REP}}^{u}+\lambda\mathcal{L}_{\text{REP}}^{s},$ (12)
@@ 4.5 Final Objective
where $\beta$ controls the contributions of ETF loss and representation learning loss.
@@
(a) Clustering with the ground-truth number of categories $K$ given
GCD (CVPR22) 51.3 56.6 48.7 39.0 57.6 29.9 45.0 41.1 46.9 35.4 51.0 27.0 73.0 76.2 66.5 74.1 89.8 66.3
DCCL (CVPR23) 63.5 60.8 64.9 43.1 55.7 36.2 - - - - - - 75.3 76.8 70.2 80.5 90.5 76.2
PromptCAL 62.9 64.4 62.1 50.2 70.1 40.6 52.2 52.2 52.3 37.0 52.0 28.9 81.2 84.2 75.3 83.1 92.7 78.3
GPC (ICCV23) 55.4 58.2 53.1 42.8 59.2 32.8 46.3 42.5 47.9 - - - 77.9 85.0 63.0 76.9 94.3 71.0
SimGCD (ICCV23) 60.3 65.6 57.7 53.8 71.9 45.0 54.2 59.1 51.8 44.0 58.0 36.4 80.1 81.2 77.8 83.0 93.1 77.9
PIM (CVPR23) 62.7 75.7 56.2 43.1 66.9 31.6 - - - 42.3 56.1 34.8 78.3 84.2 66.5 83.1 95.3 77.0
TRAILER (CVPR24) 65.1 71.3 61.9 55.4 71.7 47.6 54.5 62.6 50.5 44.5 57.0 37.8 - - - - - -
SelEx (ECCV24) 73.6 75.3 72.8 58.5 75.6 50.3 57.1 64.7 53.3 - - - 82.3 85.3 76.3 83.1 93.6 77.8
CMS (CVPR24) 68.2 76.5 64.0 56.9 76.1 47.6 56.0 63.4 52.3 36.4 54.9 26.4 82.3 85.7 75.5 84.7 95.6 79.2
SPT (ICLR24) 65.8 68.8 65.1 59.0 79.2 49.3 59.3 61.8 58.1 43.4 58.7 35.2 81.3 84.3 75.6 85.4 93.2 81.4
@@
(b) Clustering without the ground-truth number of categories $K$ given
GCD (CVPR22) 51.1 56.4 48.4 39.1 58.6 29.7 - - - 37.2 51.7 29.4 70.8 77.6 57.0 77.9 91.1 71.3
GPC (ICCV23) 52.0 55.5 47.5 38.2 58.9 27.4 43.3 40.7 44.8 36.5 51.7 27.9 75.4 84.6 60.1 75.3 93.4 66.7
PIM (CVPR23) 62.0 75.7 55.1 42.4 65.3 31.3 - - - 42.0 55.5 34.7 75.6 81.6 63.6 83.0 95.3 76.9
CMS (CVPR24) 64.4 68.2 62.2 51.7 68.9 43.4 55.2 60.6 52.4 37.4 56.5 27.1 79.6 83.2 72.3 81.3 95.6 74.2
@@ 5.1 Experimental Setup
Datasets. Our evaluation spans six image classification benchmarks, including both generic and fine-grained datasets. For generic datasets, we use CIFAR-100 krizhevsky2009learning and ImageNet-100 deng2009imagenet ; tian2020contrastive . For fine-grained datasets, we evaluate on CUB-200 wah2011caltech , Stanford Cars krause20133d , FGVC Aircraft maji2013fine , and Herbarium19 tan2019herbarium . To separate categories into known and novel categories, we follow the SSB split protocol li2020ssb for the fine-grained datasets. For CIFAR-100 and ImageNet-100, we perform a random category split using a fixed seed, consistent with previous studies.
Implementation Details. Our method utilizes the pre-trained DINO ViT-B/16 caron2021emerging ; dosovitskiy2021image as the backbone network. We fine-tune the final layer of the image encoder along with a projection head, following prior work GCD2022 ; CMS2024 . The projection head consists of an MLP with a 768-dimensional input, a 2048-dimensional hidden layer, and a 768-dimensional output, followed by a GeLU activation function hendrycks2016gaussian . The learning rate is set to 0.1. Other hyperparameters, including batch size, temperature $\tau_{s}$, weight decay $\lambda$, and the number of augmentations, are set to 128, 0.07, $1e^{-4}$, and 2, respectively, in accordance with previous studies. All experiments are conducted on an NVIDIA 3090 GPU. Further implementation details can be found in the Appendix.
5.2 Comparison with State of the Art Methods
We evaluate our NC-GCD framework on several fine-grained and generic classification benchmarks to demonstrate its efficacy in GCD. Table 1 presents a comparison between our method and the previous methods GCD2022 ; DCCL2023 ; PromptCAL2023 ; GPC2023 ; SimGCD2023 ; PIM2023 ; CMS2024 ; SPT2024 ; SelEx2024 ; Xiao2024TRA , evaluated on both fine-grained and generic datasets with and without access to the ground-truth (GT) number of categories K for clustering. For the case without access to GT K, we estimate the K through clustering, which is then used to align the ETF prototypes. Meanwhile, to comprehensively evaluate the performance of various methods across Fine-Grained Datasets, Generic Classification Datasets, and All datasets, Table 2 presents the average performance.
Fine-Grained Datasets. On fine-grained benchmarks such as CUB-200, Stanford Cars, FGVC Aircraft, and Herbarium19, NC-GCD outperforms previous methods, whether or not the ground-truth K is provided. As shown in Table 2a, when the GT K is available, NC-GCD achieves an average all-category accuracy of 60.3%, surpassing SPT by 3.4% and improving novel category accuracy over SPT by 4.8%. Even without GT K, as shown in Table 2b, NC-GCD remains robust, surpassing CMS by 3.3% in all-category accuracy. Notably, on CUB-200, NC-GCD outperforms CMS by 5.9% in all-category accuracy and 7.2% in novel category accuracy, demonstrating strong generalization without the number of categories. Fine-grained datasets present higher category confusion due to minimal intra-category variation, and our method effectively mitigates this and enhances category separability.
Generic Classification Datasets. NC-GCD achieves strong performance on standard classification datasets such as CIFAR-100 and ImageNet-100, as shown in Table 2. With known K, NC-GCD attains an all-category accuracy of 82.7% on CIFAR-100 and 88.4% on ImageNet-100, outperforming CMS and SPT. Notably, novel category accuracy improves by 1.7% on CIFAR-100 and 4.1% on ImageNet-100. When K must be estimated, NC-GCD maintains its competitive edge, achieving 85.7% all-category accuracy on ImageNet-100, surpassing CMS by 4.4%, demonstrating its adaptability in generic classification scenarios.
Across both fine-grained and generic datasets, NC-GCD consistently achieves well performance. As shown in Table 2a, with GT K, NC-GCD reaches the highest all-category accuracy of 68.7%, surpassing SPT by 3.0% and improving novel category accuracy by 4.1%. Even without GT K, NC-GCD remains robust, surpassing CMS by 3.1% in all-category accuracy and 4.4% in novel category accuracy. By pre-assigning ETF prototypes, NC-GCD effectively improves category separability, mitigating category confusion and ensuring balanced learning across old and novel categories. These results highlight the robustness and versatility of NC-GCD for real-world GCD applications.
Method Fine-grained Avg Classification Avg All Avg
@@ 5.2 Comparison with State of the Art Methods
(a) Clustering with the ground-truth number of categories K given
@@
Ours (NC-GCD) 60.3 67.6 56.7 85.5 89.8 81.4 68.7 75.0 64.9
Improv. over SPT +3.4 +0.5 +4.8 +2.1 +1.0 +2.9 +3.0 +0.7 +4.1
(b) Clustering without the ground-truth number of categories K given
@@
Ours (NC-GCD) 55.5 64.6 50.9 83.1 89.8 77.3 64.7 73.0 59.7
Improv. over CMS +3.3 +1.0 +4.6 +2.6 +0.4 +4.0 +3.1 +0.8 +4.4
@@ 5.3 Ablation Study
Unsupervised ETF Alignment: This component yields the most significant improvements, especially in fine-grained datasets. As shown in Table 3, enabling unsupervised ETF alignment alone leads to substantial gains in novel category accuracy. The model effectively learns robust representations, particularly for novel categories, where differentiation between visually similar categories is crucial.
Supervised ETF Alignment: This component stabilizes and preserves old categories’ accuracy. By aligning labeled features with their corresponding ETF prototypes, it ensures the retention of learned knowledge while maintaining high performance on known categories. Though its contribution to novel category accuracy is smaller than that of unsupervised alignment, it plays a crucial role in preventing forgetting and maintaining model stability.
Synergistic Combination: The full model, integrating both alignments, achieves the highest overall accuracy, with an average all-category accuracy of 68.7%, significantly surpassing the baseline. This combination enhances novel category discovery while preserving old category performance. Specifically, the average novel category accuracy increases by 7.4%, and the overall accuracy improves by 5.4%. The results highlight the synergy between supervised and unsupervised alignment, demonstrating the NC-GCD effectiveness in balancing learning across known and novel categories.
Table 3: Performance comparison with Supervised and Unsupervised ETF Alignment.
Sup ETF Alignment Unsup ETF Alignment CUB Herbarium 19 ImageNet100 Six Datasets Avg
All Old New All Old New All Old New All Old New
67.7 75.7 63.9 36.5 55.0 26.5 84.4 94.0 80.8 63.3 73.7 57.5
69.6 75.8 66.5 37.6 56.7 28.1 85.0 95.3 79.3 64.6 75.2 58.3
75.7 71.3 77.8 47.2 60.0 40.3 87.6 94.5 84.1 68.0 73.2 64.9
74.8 76.8 73.8 46.4 58.4 40.7 88.4 94.1 85.5 68.7 74.9 64.9
Improv. over baseline +7.1 +1.1 +9.9 +9.9 +3.4 +14.2 +4.0 +0.1 +4.7 +5.4 +1.2 +7.4
Table 4: Ablation study of our SCM module.
SCM CUB Herbarium 19 ImageNet100
All Old New All Old New All Old New
× 70.3 71.0 70.0 42.6 54.3 36.3 84.3 88.9 82.0
75.7 71.3 77.8 47.2 60.0 40.3 87.6 94.5 84.1
Improv. +5.4 +0.3 +7.8 +4.6 +5.7 +4.0 +3.3 +5.6 +2.1
The Effectiveness of SCM. To evaluate the impact of the Semantic Consistency Matcher (SCM), we conduct ablation studies on CUB-200, ImageNet100, and Herbarium 19. As shown in Table 5.3, enabling SCM consistently improves performance on all three datasets. On Herbarium 19, which poses greater challenges due to its large number of categories and high intra-category variance, SCM provides more significant improvements, with gains of 4.6% for all categories and 5.7% for novel categories, demonstrating its effectiveness in complex classification tasks. By mitigating clustering fluctuations, SCM ensures stable pseudo-label assignments. These results highlight its crucial role in stabilizing training, particularly in datasets with large category numbers.
Figure 3: Ablation study of the $\alpha$ on CUB-200.
The Influence of the ETF Scope Coefficient $\alpha$. The ETF Scope Coefficient $\alpha$ plays a crucial role in determining the number of samples selected for alignment with the ETF prototype. By controlling the percentage of high-confidence samples, $\alpha$ influences the model’s ability to focus on either old or novel categories. As seen in Fig. 3, when $\alpha$ is too small, the model may not leverage enough high-confidence samples, leading to suboptimal performance. On the other hand, as $\alpha$ increases, novel category accuracy improves significantly, especially in fine-grained datasets where category distinctions are subtle. However, when $\alpha$ exceeds a certain threshold, the model’s focus on new categories comes at the expense of old category accuracy, highlighting the trade-off between the two. The best balance is achieved at $\alpha=0.8$, where both old and novel category accuracy reach optimal values, demonstrating the importance of selecting an appropriate $\alpha$.